HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

ANALYSIS OF HIGH-RESOLUTION SPECTRA OF SiF_{4} COMBINATION BANDS

\underline{M. A. MERKULOVA} , National Research Tomsk Polytechnic University, 30, av. Lenina, 634050 Tomsk, Russia; \underline{V. BOUDON}, ICB, 9, av. Alain Savary, B.P. 47870, F-21078 Dijon, France; \underline{L. MANCERON} , Synchrotron SOLEIL, AILES Beamline,

L’Orme des Merisiers, BP48, F-91192 St-Aubin Cedex, France

In this study, the spectrum of the silicon tetrafluoride molecule (SiF_{4}) was examined in five different combination band ranges. The energy structure of the vibrational states, specifically (v_{1} = v_{3} = 1), (v_{1} = v_{4} = 1), (v_{2} = v_{3} = 1), (v_{2} = v_{4} = 1), and very weak (v_{3} = v_{4} = 1), was determined. More than 8120 transitions were assigned with maximum rotational quantum number J_{max} values reaching up to 60 (up to 82 for ν_{1}+ν_{3}).

These results allowed for the precise determination of molecular parameters, with root mean square deviations for line positions as low as a few 10^{−4} cm^{−1}.

In the case of the ν_{1}+ν_{3} band, the ^{29}SiF_{4} and ^{30}SiF_{4} isotopologues could also be assigned and fitted (figure below). These results will allow the calculation of hot bands like ν_{3}+ν_{2}-ν_{2}, etc., in the regions of strong atmospheric absorption.

The obtained results hold significant value for their future application in the precise semi-empirical determination of the intramolecular potential function of tetrafluorosilane. Furthermore, they are of critical importance for the subsequent analysis of the line intensities associated with hot bands of this molecule. Part of this work was recently published^{1}.

1M. Merkulova, V. Boudon and L. Manceron, J. Mol. Spectrosc. 391, 111738 (2023).

p-number: p095  Submitted on Tue, 13 Jun 2023 14:17:48 +0200

HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

Overview of the ν_{1}+ν_{3} spectrum, compared to the simulation for all isotopologues. The

insert details a part of the Q branch region.

p-number: p095  Submitted on Tue, 13 Jun 2023 14:17:48 +0200